# Optimizing Spin Dressing Sensitivity for the nEDMSF Experiment

### v1

@@ 2 Square-Wave-Modulated Critical Dressing
rapidly enough that the precession about the pseudomagnetic field can be ignored [2]. Here we have used $\widehat{EB}\equiv\hat{E}\cdot\hat{B}$, refer to $\phi_{d}$ as the ”dressing angle”, and
@@
is the frequency shift resulting from the combination of the true $d_{n}$, a false $d_{n}$ if one has been injected into the datastream to blind the analysis [9], the gravitational offset between the UCN and the 3He comagnetometer [10], and the combination of perturbative magnetic fields arising from residual magnetic field gradients and the applied electric field [10, 11, 12, 13, 14, 15, 16]. We refer to this approach as ”square-wave-modulated critical dressing”.

### v2

@@ 2 Square-Wave-Modulated Critical Dressing
rapidly enough that the precession about the pseudomagnetic field can be ignored [2]. Here we have used $\widehat{EB}\equiv\hat{E}\cdot\hat{B}$ and refer to $\phi_{d}$ as the ”dressing angle”.
@@
is the frequency shift resulting from the combination of the true $d_{n}$, a false $d_{n}$ if one has been injected into the datastream to blind the analysis [9], the gravitational offset between the UCN and the 3He comagnetometer [10], the combination of perturbative magnetic fields arising from residual magnetic field gradients and the applied electric field [10, 11, 12, 13, 14, 15, 16], and any static offset between $\gamma_{n}^{\prime}$ and $\gamma_{3}^{\prime}$. We refer to this approach as ”square-wave-modulated critical dressing”.